OpenStudy (anonymous):

How would i write an equation of a line in point-slope form with a slope of -3/4 and passing through the point (1, -5).

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -3/4 (1,-5)...y1 = 1 and x1 = -5 now we just sub y - 1 = -3/4(x - (-5) = y - 1 = -3/4(x + 5) understand ? you do nothing else if you want it in point slope form...you do not distribute the -3/4 through the parenthesis

OpenStudy (anonymous):

wouldn't the y be -5 and the x be 1?

OpenStudy (texaschic101):

oops.. (1,-5).....x1 = 1 and y1 = -5 y - (-5) = -3/4(x - 1) = y + 5 = -3/4(x - 1).....sorry about that

OpenStudy (anonymous):

no problem:] I def understand how to do this now!

OpenStudy (texaschic101):

lol.....I am happy to help..or at least try to help

OpenStudy (anonymous):

thanks for your help!